If kaylee rented a bike front sarawongs bike shop it cost $10 plus $5 per hour if kaylee paid $20 how many hours ahead rent the bike for. write an equation
Answer by stanbon(75887)

If kaylee rented a bike from sara wongs bike shop it cost $10 plus $5 per hour. If kaylee paid $20, how many hours did she rent the bike for.
Write an equation.
----
Cost = C(x) = 10 + 5x, where x is the number of hours.
--------------
Solve:
20 = 10 + 5x
5x = 10
x = 2 hrs.
-----
Cheers,
Stan H.
-----------